# Dark matter admixed neutron stars with a realistic nuclear equation of state from chiral nuclear interactions

### v1

@@ IV DANS properties
Table 1: Values for the gravitational mass of ordinary matter and dark matter and total fraction of dark matter obtained for $m_{\chi}=1\ \mathrm{GeV}$, $\rho_{cOM}=2\times 10^{15}\ \mathrm{g/cm^{3}}$ and different values of the dark matter central mass-density.
Figure 2: Results for equilibrium configurations of DANS in the $(\rho_{cOM},\rho_{cDM})$ plane for the case $m_{\chi}=1\ \mathrm{GeV}$. The left panel shows the total gravitational mass of the star $M_{tot}=M_{tot}(\rho_{cOM},\rho_{cDM})$ for different choices of the central mass densities. The central and right panel show the quantity $\Delta R=R_{OM}-R_{DM}$ and the stellar DM fraction in percent $f_{\chi}\,[\%]$ respectively as function of the central mass densities. On each of the three panels we report the contour lines for $M_{tot}$ (left), $\Delta R$ (central) and $f_{\chi}\,[\%]$ (right). The green and blue lines in each panel delimit the region in the central mass density plane where the total gravitational mass $M_{tot}$ is compatible with the measurement of the mass of the neutron star associated with the pulsar PSR J0348+0432 ($2.01\pm 0.04\ \mathrm{M_{\odot}}$) Antoniadis et al. (2013).
Figure 3: Same as in Fig (2) but now for the the case $m_{\chi}=0.4\ \mathrm{GeV}$.
As an illustrative example, we display in Figure (1) the radial mass density profiles $\rho_{OM}(r)$ for OM (red curves) and $\rho_{DM}(r)$ for DM (black curves) obtained for a fixed value $\rho_{cOM}=2\times 10^{15}\ \mathrm{g/cm^{3}}$ of the OM fluid central density and using different values $\rho_{cDM}$ of the DM fluid central density. Each line-style in Figure (1) corresponds to the profiles obtained for a different value of $\rho_{cDM}$ (see the figure caption for more details). As a baseline, we also plot in the same figure the mass density profile of an ordinary neutron star (green line). Taking the same stellar configurations considered in Figure (1), we report in Table (1) the total gravitational masses (eq.(4)) of the OM and DM stellar components, the corresponding radii $R_{OM}$ and $R_{DM}$ and the stellar DM fraction $f_{\chi}$ defined as
$f_{\chi}=\frac{M_{DM}}{M_{tot}}.$ (6)
@@
The results of our systematic calculations for hydrostatic equilibrium configurations of DANS in the $(\rho_{cOM},\rho_{cDM})$ plane are reported in Fig (2) for the case $m_{\chi}=1\ \mathrm{GeV}$. In particular, in the left panel we report, using a color scale, the total gravitational mass $M_{tot}$ of the star. The black dotted lines represent the contour lines ($M_{tot}=\mathrm{const}$), whereas the green and blue lines (reported in each of the three panels) are the contour lines which mark the region in the central mass densities plane where the calculated $M_{tot}$ is compatible with the measured mass ($2.01\pm 0.04\ \mathrm{M_{\odot}}$) Antoniadis et al. (2013) of the neutron star associated to the pulsar PSR J0348+0432. In the central panel of Fig (2), we report, using a color scale, the difference $\Delta R=R_{OM}-R_{DM}$ between the radii of the OM and DM fluid distributions. The white dashed line labeled by ”0.0” represents the contour line $\Delta R=0$ and thus marks the boundary between DANS with a DM core (on the right of the $\Delta R=0$ line) and those with a DM halo (on the left of the $\Delta R=0$ line). Finally, in the right panel of Fig (2) we plot the DM fraction (in percent) $f_{\chi}\,[\%]$ and the contour lines for this quantity (red dash-dotted lines). The results in the right panel of Fig (2) ($m_{\chi}=1\,\mathrm{GeV}$) clearly show that only DANS with a low DM fraction ($f_{\chi}<4\,\%$) are compatible with the measured mass of the neutron star associated with PSR J0348+0432 and that these configurations are all characterized by the presence of a DM core (see middle panel). In Fig. (3) we show our results for DANS properties in the case $m_{\chi}=0.4\ \mathrm{GeV}$. Now, in addition to the region of the $(\rho_{cOM},\rho_{cDM})$ plane populated by DANS with a DM core and $f_{\chi}\lesssim 4.6\,\%$ and having $M_{tot}\sim 2\,\mathrm{M_{\odot}}$, there is a second region of the $(\rho_{cOM},\rho_{cDM})$ plane populated by DANS with $M_{tot}\gtrsim 2\,\mathrm{M_{\odot}}$. All members of this second group of DANS have an extended DM halo and are almost completely formed by DM. In the $m_{\chi}=0.4\ \mathrm{GeV}$ case a large region of the $(\rho_{cOM},\rho_{cDM})$ plane is populated by DANS with an extended DM halo and having a high DM fraction.
@@
IV.1 Mass-radius relation for fixed $f_{\chi}$
@@
In figure (4), for the case $m_{\chi}=1\ \mathrm{GeV}$, we show the total gravitational mass of the star $M_{tot}$ as a function of $\rho_{cOM}$ (left panel) and $M_{tot}$ as a function of the radius $R_{OM}$ of the OM fluid distribution (right panel). The surface of the OM fluid distribution is in fact the place where the electromagnetic radiation, that allows to determine $R_{OM}$, is generated. The light blue dash-dotted curve in both panels represents the ordinary neutron star sequence ($f_{\chi}=0$). The other two curves represent DANS sequences with $f_{\chi}=3\%$ (orange dash-dotted lines) and $f_{\chi}=6\%$ (green dash-dotted lines). These two stellar sequences represent DANS having a DM core (see central and right panels in Fig (2)). The light green horizontal band represents the measured mass $(2.08\pm 0.07)\ \mathrm{M_{\odot}}$ Fonseca et al. (2021) of the neutron star associated to the pulsar PSR J0740+6620, whereas the pink horizontal band represents the measured mass $(2.01\pm 0.04)\ \mathrm{M_{\odot}}$ Antoniadis et al. (2013) of the neutron star associated to the pulsar PSR J0348+0432. As we can see from the results reported in figure (4), the presence of DM with $m_{\chi}=1\ \mathrm{GeV}$ reduces the value of the maximum gravitational mass $M_{max}\equiv M_{tot,max}$ of the star with respect to the one for ordinary ($f_{\chi}=0$) neutron stars. Considering DANS sequences with $f_{\chi}\gtrsim 3\%$ the calculated maximum mass is no longer compatible with the present measured masses of ”heavy” neutron stars with $M\sim 2\,\mathrm{M_{\odot}}$ Fonseca et al. (2021); Antoniadis et al. (2013). We also note that the $M_{tot}(\rho_{cOM})$ curve is shifted towards higher value of the ordinary matter central density $\rho_{cOM}$ compared to the ordinary neutron stars case ($f_{\chi}=0\%$) as $f_{\chi}$ increases. The presence of DM also affects the value of the OM radius $R_{OM}$ (right panel) making the OM distribution more compact Ellis et al. (2018); Ivanytskyi et al. (2020). For example, considering an ordinary neutron star with a gravitational mass $M=1.4\ \mathrm{M_{\odot}}$, a DM fraction $f_{\chi}=6\%$ reduces the stellar observable radius $R_{OM}$ by $\sim 0.6\ \mathrm{km}$ i.e by $\sim 5\,\%$. These changes in the stellar radius are within the bounds of uncertainty of present instruments, as NICER, for measuring neutron star radii. The orange region, in the right panel of Fig. (4), marks the NICER measurement for the mass and radius of the neutron star in PSR J0030+0451 Miller et al. (2019) while the gray and blue regions represent the portion of the mass-radius plane compatible with the data obtained from the GW170817 event (at $90\%\,CL$) Abbott et al. (2018).
@@
Figure 4: Total gravitational mass of the star $M_{tot}$ as a function of $\rho_{cOM}$ (left panel) and $M_{tot}$ as a function of the radius $R_{OM}$ of the OM fluid distribution (right panel) for different values of the DM fraction ($f_{\chi}=0\,\%,3\,\%,6\,\%$) and in the case $m_{\chi}=1\ \mathrm{GeV}$. The red star symbol on the top of each curve marks the maximum mass configuration. The light green and the pink horizontal bands represent respectively the measured mass of the neutron star associated to the pulsar PSR J0740+6620 Fonseca et al. (2021) and to PSR J0348+0432 Antoniadis et al. (2013). In the right panel, the orange region marks the NICER measurement for the mass and radius of the neutron star in PSR J0030+0451 Miller et al. (2019) while the gray and blue regions represent the portion of the mass-radius plane compatible with the data obtained from the GW170817 event (at $90\%\,CL$) Abbott et al. (2018).
@@
In Fig. (5) we report DANS sequences for DM particles having a mass $m_{\chi}=0.1\ \mathrm{GeV}$ and for fixed DM fraction. In this case we get stars with a DM halo. This leads to a growth of the total gravitational maximum mass $M_{max}$ of the star as $f_{\chi}$ increases. Since the ordinary neutron star sequence, calculated with the BL EOS, is compatible with the present measured masses of ”heavy” neutron stars, the agreement with the same observational data will clearly hold for DANS sequences in the $m_{\chi}=0.1\ \mathrm{GeV}$ case.
Figure 5: Same as in Fig (4) but now for the the case $m_{\chi}=0.1\ \mathrm{GeV}$ and DM fraction $f_{\chi}=0\,\%,4\,\%,10\,\%$.
@@
To investigate in details how the maximum mass $M_{max}$ of DANS depends on the DM fraction $f_{\chi}$ and on the DM particle mass $m_{\chi}$, we show in Fig. (6) our results for the function $M_{max}(m_{\chi},f_{\chi})$ calculated for a few constant values of $m_{\chi}$ (curves in different colors). The black heavy dot on top of each curve denotes the minimum of the function $M_{max}(f_{\chi})$. We see that for $m_{\chi}\approx 0.188\ \mathrm{GeV}$ the minimum of the orange curve is equal to $2\ \mathrm{M_{\odot}}$ (a value we take as a reference for the measured masses of “heavy” neutron stars): this result means that for $m_{\chi}\leq 0.188\ \mathrm{GeV}$ we obtain a value for $M_{max}$ that is always compatible with measured neutron star masses for any value of $f_{\chi}$.
@@
Following Ivanytskyi et al. (2020); Sagun et al. , for any fixed value of $m_{\chi}\geq 0.188\ \mathrm{GeV}$, we define a critical value of the DM fraction, and denote it with $f_{\chi}^{crit}=f_{\chi}^{crit}(m_{\chi})$ , as the solution of the equation $M_{max}(f_{\chi},m_{\chi})=2\ \mathrm{M_{\odot}}$.
From figure (6) we see that for $0.188\ \mathrm{GeV}\leq m_{\chi}\leq 0.4\ \mathrm{GeV}$ we get two critical values for $f_{\chi}$ and we observe that the largest of the two roots tends to $f_{\chi}=100\%$ as $m_{\chi}$ increases. For the considered values $m_{\chi}\geq 0.4\ \mathrm{GeV}$ we get one critical value for $f_{\chi}$.
The exact value of $m_{\chi}>0.188\ \mathrm{GeV}$ that marks the transition from two critical values to one critical value for $f_{\chi}$ must lie in the range $0.560.4<m_{\chi}/\mathrm{GeV}<$ 0.56. For this particular value of $m_{\chi}$, the largest root of the equation $M_{lim}(f_{\chi},m_{\chi})=2\ \mathrm{M_{\odot}}$ is $f_{\chi}\approx 100\%$.
The critical DM fraction is shown in Fig. (7) by the heavy black curve labeled $f_{\chi}^{crit}(m_{\chi})$. The green region in the same figure thus represents the portion of the $(m_{\chi},f_{\chi})$ plane where the calculated maximum mass for DANS is compatible with the present measured masses of ”heavy” neutron stars with $M_{tot}\sim 2\,\mathrm{M_{\odot}}$ Fonseca et al. (2021); Antoniadis et al. (2013). It is useful to observe that, in the range $10\leq m_{\chi}/\mathrm{GeV}\leq 100$ the critical DM curve $f_{\chi}^{crit}(m_{\chi})$ is well described by the following equation
@@
Figure 7: Critical value of $f_{\chi}$ as a function of $m_{\chi}$. In the green region and along the black line the maximum mass for DANS is compatible with observation.
IV.2 Proton fraction and direct URCA processes in DANS
As we show in this subsection, the presence of DM influences not only the bulk properties (mass, radius, maximum mass) of a neutron star, but in addition has indirect sizeable effect (through gravitational interaction) on the particle fractions of the OM fluid constituents when considering DANS with a fixed total gravitational mass. In particular the value of $f_{\chi}$ influences the proton fraction $x_{p}=n_{p}/n$ (with $n=n_{n}+n_{p}$ being the baryon number density) in $\beta$-stable nuclear matter. As has been shown in Ref. Lattimer et al. (1991), when the proton fraction $x_{p}$ is larger than a threshold value, denoted as $x_{p}^{durca}$, the so-called direct URCA processes $n\rightarrow p+e^{-}+\bar{\nu}_{e}\,$, ${}p+e^{-}\rightarrow n+\nu_{e}\,,$ can occur in neutron star matter. The direct URCA processes enhances neutrino emission and neutron star cooling rates by a very large factor compared to the so-called modified URCA processes $(n,p)+n\rightarrow p+e^{-}+\bar{\nu}_{e}\,$, ${}(n,p)+p+e^{-}\rightarrow n+\nu_{e}\,,$ causing a fast cooling of the star Lattimer et al. (1991); Page et al. (2004).
In $\beta$-stable nuclear matter the threshold proton fraction for direct URCA processes can be written Bombaci and Logoteta (2018)

### v2

@@ 4 DANS properties
Table 1: Values for the gravitational mass of ordinary matter and dark matter and total fraction of dark matter obtained for $m_{\chi}=1\ \mathrm{GeV}$, $\rho_{cOM}=1.5\times 10^{15}\ \mathrm{g/cm^{3}}$ and different values of the dark matter central mass-density.
Figure 2: Results for equilibrium configurations of DANS in the $(\rho_{cOM},\rho_{cDM})$ plane for the case $m_{\chi}=1\ \mathrm{GeV}$. The left panel shows the total gravitational mass of the star $M_{tot}=M_{tot}(\rho_{cOM},\rho_{cDM})$ for different choices of the central mass densities. The central and right panel show the quantity $\Delta R=R_{OM}-R_{DM}$ and the stellar DM fraction in percent $f_{\chi}\,[\%]$ respectively as function of the central mass densities. On each of the three panels we report the contour lines for $M_{tot}$ (left), $\Delta R$ (central) and $f_{\chi}\,[\%]$ (right). The green and blue lines in each panel delimit the region in the central mass density plane where the total gravitational mass $M_{tot}$ is compatible with the measurement of the mass of the neutron star associated with the pulsar PSR J0348+0432 ($2.01\pm 0.04\ \mathrm{M_{\odot}}$) (Antoniadis et al., 2013).
Figure 3: Same as in Fig (2) but now for the case $m_{\chi}=0.4\ \mathrm{GeV}$.
As an illustrative example, we display in Figure (1) the radial mass density profiles $\rho_{OM}(r)$ for OM (red curves) and $\rho_{DM}(r)$ for DM (black curves) obtained for a fixed value $\rho_{cOM}=1.5\times 10^{15}\ \mathrm{g/cm^{3}}$ of the OM fluid central density and using different values $\rho_{cDM}$ of the DM fluid central density. Each line-style in Figure (1) corresponds to the profiles obtained for a different value of $\rho_{cDM}$ (see the figure caption for more details). As a baseline, we also plot in the same figure the mass density profile of an ordinary neutron star (green line). Taking the same stellar configurations considered in Figure (1), we report in Table (1) the total gravitational masses (eq.(4)) of the OM and DM stellar components, the corresponding radii $R_{OM}$ and $R_{DM}$ and the stellar DM fraction $f_{\chi}$ defined as
$f_{\chi}=\frac{M_{DM}}{M_{tot}}$ (6)
@@
The results of our systematic calculations for hydrostatic equilibrium configurations of DANS in the $(\rho_{cOM},\rho_{cDM})$ plane are reported in Fig. (2) for the case $m_{\chi}=1\ \mathrm{GeV}$. In particular, in the left panel we report, using a color scale, the total gravitational mass $M_{tot}$ of the star. The black dotted lines represent the contour lines ($M_{tot}=\mathrm{const}$), whereas the green and blue lines (reported in each of the three panels) are the contour lines which mark the region in the central mass densities plane where the calculated $M_{tot}$ is compatible with the measured mass ($2.01\pm 0.04\ \mathrm{M_{\odot}}$) (Antoniadis et al., 2013) of the neutron star associated to the pulsar PSR J0348+0432. In the central panel of Fig (2), we report, using a color scale, the difference $\Delta R=R_{OM}-R_{DM}$ between the radii of the OM and DM fluid distributions. The white dashed line labeled by ”0.0” represents the contour line $\Delta R=0$ and thus marks the boundary between DANS with a DM core (on the right of the $\Delta R=0$ line) and those with a DM halo (on the left of the $\Delta R=0$ line). Finally, in the right panel of Fig (2) we plot the DM fraction (in percent) $f_{\chi}\,[\%]$ and the contour lines for this quantity (red dash-dotted lines). The results in the right panel of Fig (2) ($m_{\chi}=1\,\mathrm{GeV}$) clearly show that only DANS with a low DM fraction ($f_{\chi}<4\,\%$) are compatible with the measured mass of the neutron star associated with PSR J0348+0432 and that these configurations are all characterized by the presence of a DM core (see middle panel).
In Fig. (3) we show our results for DANS properties in the case $m_{\chi}=0.4\ \mathrm{GeV}$. Now, in addition to the region of the $(\rho_{cOM},\rho_{cDM})$ plane populated by DANS with a DM core, with a low DM fraction ($f_{\chi}\lesssim 3.3\,\%$) and having $M_{tot}\sim 2\,\mathrm{M_{\odot}}$, there is a second region of the $(\rho_{cOM},\rho_{cDM})$ plane populated by DANS with $M_{tot}\gtrsim 2\,\mathrm{M_{\odot}}$. All members of this second group of DANS have an extended DM halo and are almost completely formed by DM.
The possibility of having DANS with a large DM fraction or compact stars made entirely by fermionic DM and the possible mechanisms to form them have been discussed in several studies (e.g. (Narain et al., 2006; Kouvaris and Nielsen, 2015; Gresham and Zurek, 2019)).
@@
4.1 Mass-radius relation for fixed DM fraction
@@
In figure (4), for the case $m_{\chi}=1\ \mathrm{GeV}$, we show the total gravitational mass of the star $M_{tot}$ as a function of $\rho_{cOM}$ (left panel) and $M_{tot}$ as a function of the radius $R_{OM}$ of the OM fluid distribution (right panel). The surface of the OM fluid distribution is in fact the place where the electromagnetic radiation, that allows to determine $R_{OM}$, is generated.
@@
The light blue dash-dotted curve in both panels represents the ordinary neutron star sequence ($f_{\chi}=0$). The other two curves represent DANS sequences with $f_{\chi}=3\%$ (orange dash-dotted lines) and $f_{\chi}=6\%$ (green dash-dotted lines). These two stellar sequences represent DANS having a DM core (see central and right panels in Fig (2)). The light green horizontal band represents the measured mass $(2.08\pm 0.07)\ \mathrm{M_{\odot}}$ (Fonseca et al., 2021) of the neutron star associated to the pulsar PSR J0740+6620, whereas the pink horizontal band represents the measured mass $(2.01\pm 0.04)\ \mathrm{M_{\odot}}$ (Antoniadis et al., 2013) of the neutron star associated to the pulsar PSR J0348+0432. As we can see from the results reported in figure (4), the presence of DM with $m_{\chi}=1\ \mathrm{GeV}$ reduces the value of the maximum gravitational mass $M_{max}\equiv M_{tot,max}$ of the star with respect to the one for ordinary ($f_{\chi}=0$) neutron stars. Considering DANS sequences with $f_{\chi}\gtrsim 3\%$ the calculated maximum mass is no longer compatible with the present measured masses of ”heavy” neutron stars with $M\sim 2\,\mathrm{M_{\odot}}$ (Fonseca et al., 2021; Antoniadis et al., 2013). We also note that the $M_{tot}(\rho_{cOM})$ curve is shifted towards higher values of the ordinary matter central density $\rho_{cOM}$ compared to the ordinary neutron stars case ($f_{\chi}=0\%$) as $f_{\chi}$ increases.
The presence of DM also affects the value of the OM radius $R_{OM}$ (right panel) making the OM distribution more compact (Ellis et al., 2018; Ivanytskyi et al., 2020). For example, considering an ordinary neutron star with a gravitational mass $M=1.4\ \mathrm{M_{\odot}}$, a DM fraction $f_{\chi}=6\%$ reduces the stellar observable radius $R_{OM}$ by $\sim 0.6\ \mathrm{km}$ i.e by $\sim 5\,\%$. These changes in the stellar radius are within the bounds of uncertainty of present instruments, as NICER, for measuring neutron star radii. The orange region, in the right panel of Fig. (4), marks the NICER measurement for the mass and radius of the neutron star in PSR J0030+0451 (Miller et al., 2019), while the plum and light gray regions in the top-right part of the figure mark the NICER measurement for the mass and radius of the neutron star in PSR J0740+6620 obtained respectively in (Raaijmakers et al., 2021) and in (Miller et al., 2021). The gray and blue regions represent the portion of the mass-radius plane compatible with the data obtained from the GW170817 event (at $90\%\,CL$) (Abbott et al., 2018).
@@
Figure 4: Total gravitational mass of the star $M_{tot}$ as a function of $\rho_{cOM}$ (left panel) and $M_{tot}$ as a function of the radius $R_{OM}$ of the OM fluid distribution (right panel) for different values of the DM fraction ($f_{\chi}=0\,\%,3\,\%,6\,\%$) and in the case $m_{\chi}=1\ \mathrm{GeV}$. The red star symbol on the top of each curve marks the maximum mass configuration. The light green and the pink horizontal bands represent respectively the measured mass of the neutron star associated to the pulsar PSR J0740+6620 (Fonseca et al., 2021) and to PSR J0348+0432 (Antoniadis et al., 2013). In the right panel, the orange region marks the NICER measurement for the mass and radius of the neutron star in PSR J0030+0451 (Miller et al., 2019) while the plum and light gray regions in the top-right part of the figure mark the NICER measurement for the mass and radius of the neutron star in PSR J0740+6620 obtained respectively in (Raaijmakers et al., 2021) and in (Miller et al., 2021). The gray and blue regions represent the portion of the mass-radius plane compatible with the data obtained from the GW170817 event (at $90\%\,CL$) (Abbott et al., 2018).
@@
In Fig. (5) we report DANS sequences for DM particles having a mass $m_{\chi}=0.1\ \mathrm{GeV}$ and for fixed DM fraction. In this case, we get stars with a DM halo. This leads to a growth of the total gravitational mass of the star as $f_{\chi}$ increases for every considered value of the central density of the ordinary matter distribution and, in particular, to a growth of the maximum gravitational mass of the $M(R)$ sequence.
Figure 5: Same as in Fig (4) but now for the case $m_{\chi}=0.1\ \mathrm{GeV}$ and DM fraction $f_{\chi}=0\,\%,4\,\%,10\,\%$.
@@
To investigate in detail how the maximum mass $M_{max}$ of DANS depends on the DM fraction $f_{\chi}$ and on the DM particle mass $m_{\chi}$, we show in Fig. (6) our results for the function $M_{max}(m_{\chi},f_{\chi})$ calculated for a few constant values of $m_{\chi}$ (curves in different colors). The black heavy dot on top of each curve denotes the minimum of the function $M_{max}(f_{\chi})$ for fixed $m_{\chi}$. We see that for $m_{\chi}\approx 0.188\ \mathrm{GeV}$ the minimum of the orange curve is equal to $2\ \mathrm{M_{\odot}}$ (a value we take as a reference for the measured masses of ’heavy’ neutron stars): this result means that for $m_{\chi}\leq 0.188\ \mathrm{GeV}$ we obtain a value for $M_{max}$ that is always compatible with the measured masses of neutron stars for any value of $f_{\chi}$.
@@
For any fixed value of $m_{\chi}\geq 0.188\ \mathrm{GeV}$, we define a critical value of the DM fraction, and denote it with $f_{\chi}^{crit}=f_{\chi}^{crit}(m_{\chi})$, as the solution of the equation $M_{max}(f_{\chi},m_{\chi})=2\ \mathrm{M_{\odot}}$ (Ivanytskyi et al., 2020). From our results in Fig. (6) we see that for $0.188\,\mathrm{GeV}<m_{\chi}\leq 0.559\,\mathrm{GeV}$ there are two solutions that satisfy equation $M_{max}(f_{\chi},m_{\chi})=2\,M_{\odot}$. Thus, for example, in the case $m_{\chi}=0.4\ \mathrm{GeV}$ (red curve in our Fig. (6)) there are two groups of DANS which satisfy the condition $M_{max}(f_{\chi},m_{\chi})\geq 2\,M_{\odot}$, i.e. have masses in agreement with present measured neutron star masses. The first of these two groups is represented by DANS with a low DM fraction ($f_{\chi}\leq 3.27\%$) and having a DM core, whereas the second group consist of DANS with a high DM fraction ($f_{\chi}\geq 64.93\%$) and having a DM halo. These results are in line with our findings in Fig. (3). The value $m_{\chi}=0.559\,\mathrm{GeV}$ marks the transition from two to one solution for the critical DM fraction $f_{\chi}^{crit}$. In fact, for this value of the DM particle mass the pure ($f_{\chi}=100\%$) DM compact star sequence has a maximum mass $M_{max}(f_{\chi}=100\%)=2.00\,M_{\odot}$ (Narain et al., 2006). Our results are thus in line and extend those from (Ivanytskyi et al., 2020), where a different ordinary matter EOS was used.
The critical DM fraction is shown in Fig. (7) by the heavy black curve labeled $f_{\chi}^{crit}(m_{\chi})$. The green region in the same figure thus represents the portion of the $(m_{\chi},f_{\chi})$ plane where the calculated maximum mass for DANS is compatible with the present measured masses of ”heavy” neutron stars with $M_{tot}\sim 2\,\mathrm{M_{\odot}}$ (Fonseca et al., 2021; Antoniadis et al., 2013). It is useful to observe that, in the range $10\leq m_{\chi}/\mathrm{GeV}\leq 100$ the critical DM curve $f_{\chi}^{crit}(m_{\chi})$ is well described by the following equation
@@
Figure 7: Critical value of $f_{\chi}$ as a function of $m_{\chi}$. In the green region and along the black line the maximum mass for DANS is compatible with observation. The vertical red dashed line represents the value $m_{\chi}=0.188\,\mathrm{GeV}$. Below that value, the maximum gravitational mass obtained for the $M_{tot}(R_{OM})$ sequence is always compatible with the measured masses of neutron stars for any value of $f_{\chi}$ (see also Fig. 6). The horizontal blue dashed line indicates a conservative estimate of the dark matter fraction that a $2\ \mathrm{M_{\odot}}$ neutron star in the Galactic central region can attain due to dark matter accretion during both the main sequence (MS) stellar phase and the neutron star phase (see Sec.5).
4.2 Proton fraction and direct URCA processes in DANS
As we show in this subsection, the presence of DM influences not only the bulk properties (mass, radius, maximum mass) of a neutron star, but in addition has indirect sizable effect (through gravitational interaction) on the particle fractions of the OM fluid constituents when considering DANS with a fixed total gravitational mass. In particular the value of $f_{\chi}$ influences the proton fraction $x_{p}=n_{p}/n$ (with $n=n_{n}+n_{p}$ being the baryon number density) in $\beta$-stable nuclear matter.
As has been shown in Ref. (Lattimer et al., 1991), when the proton fraction $x_{p}$ is larger than a threshold value, denoted as $x_{p}^{durca}$, the so-called direct URCA processes $n\rightarrow p+e^{-}+\bar{\nu}_{e}\,$, ${}p+e^{-}\rightarrow n+\nu_{e}\,,$ can occur in neutron star matter. The direct URCA processes enhance neutrino emission and neutron star cooling rates by a very large factor compared to the so-called modified URCA processes $(n,p)+n\rightarrow p+e^{-}+\bar{\nu}_{e}\,$, ${}(n,p)+p+e^{-}\rightarrow n+\nu_{e}\,,$ causing a fast cooling of the star (Lattimer et al., 1991; Page et al., 2004).
In $\beta$-stable nuclear matter the threshold proton fraction for direct URCA processes can be written (Bombaci and Logoteta, 2018)